# MMiC: Mitigating Modality Incompleteness in Clustered Federated Learning

### v1

@@ 4.4 Variants of MMiC
5 Experiments
Table 2: Global and personalized performance compared with different algorithms. Bold with * indicated the best performance, while bold black indicated excellent or stable top three performance. Since the PmcmFL algorithm is only for classification tasks, it can’t be compared in Flick30k. MR stands for Modality Reconstruction and GF is GMC-Fusion.
@@ 5 Experiments
5.1 Datasets, Models and Metrics
Our evaluation of MMiC is based on two multimodal classification datasets and one multimodal retrieval dataset. Because each dataset has distinct characteristics, we followed (Feng et al., 2023; Jeni et al., 2013) in employing different base models and evaluation metrics for each dataset. We also provide learning configurations in Appendix A. CrisisMMD (Ofli et al., 2020). It is a multi-modal dataset for classification consisting of 18,126 samples annotated tweets and images. We brought in a lightweight multi-modal model that used pre-trained embedding layers for input and concatenates shared representations (Gupta et al., 2024; Feng et al., 2023). Following (Feng et al., 2023), we employed the F1 (Goutte & Gaussier, 2005) score to evaluate the performance of CrisisMMD.
UPMC-Food101 (Gallo et al., 2020). This dataset was sourced from Food101, containing 101,000 image-text pairs across 101 food categories. We used CLIP with the official ViT version of the backbone pre-trained model (Radford et al., 2021). We froze the first 11 layers of attention while retaining the last layer of attention and the final output linear layer for parameter updates. We employed accuracy to evaluate its performance.
@@ 5.1 Datasets, Models and Metrics
5.2 Benchmark Algorithms
We compared MMiC with selected representative works and state-of-the-art works.
FedAvg (McMahan et al., 2017) represented a basic federated architecture where multiple clients collaborated with a central server. FedOpt (Reddi et al., 2021a) proposed three specialized federated aggregation methods, building upon FedAvg. FedSoft (Ruan & Joe-Wong, 2022), as a soft clustering method, introduced a technique that permitted a client to belong to multiple clusters. PACFL (Vahidian et al., 2023), an advanced CFL methodology, employed Singular Value Decomposition (SVD) to deconstruct the dataset into its constituent subspaces, subsequently applying hierarchical clustering techniques to these subspaces for effective clustering. CreamFL (Yu et al., 2023), was a MFL algorithm that employed inter-modal contrastive learning to address the issue of modality imbalance. PmcmFL (Bao et al., 2024), was another MFL for addressing missing modality. They utilized prototype aggregation to address the issue of missing modalities in classification tasks.
5.3 Settings
@@
Cluster and Clients Settings. In this experiment, we employed two distance metrics for clustering: Locality Sensitive Hashing (LSH) sketch (Yang et al., 2024; Liu et al., 2023) and SVD decomposition (Vahidian et al., 2023). All experiments used the LSH sketch for clustering unless otherwise noted. Each cluster had a test set visible to all clients within the cluster. Additionally, the global server provided accessibility to the test set for all clients. We simulated the Non-IID scenario by assigning a label probability table to each client utilizing the Dirichlet distribution (Teh, 2010) with $\beta_{diri}=0.5$. For all the experiments, we set the number of clients to 50 so that ensured each client had at least 200 training data. Our client selection strategy comprised an initial phase of 100 rounds with random selection, followed by a subsequent phase of 200 rounds utilizing Banzhaf client selection (300 global rounds).
@@
Missing Modality Settings. Our setting with missing modalities was governed by two variables: clients with missing modality (mc) and missing modality rate (mm), denoting the proportion of all clients selected to have missing modalities and the proportion of missing samples on these clients, respectively. Our main experiment (Section 5.4) offered three levels of modality incompleteness settings: (mm@0.2, mc@0.3), (mm@0.2, mc@0.5) and (mm@0.5, mc@0.5). each client has a fixed number of missing samples, determined by the mm rate. In each round, among the clients selected for aggregation, we use mc to decide whether the selected clients will activate missing data. Since the selected clients use a random training subset for training in each round, the maximum missing modality rate for a client in that round is $min(\frac{missing\ samples}{subset\ samples},1)$.
5.4 Overall Performance
We evaluated MMiC based on global performance (gbp) and personalized performance (plp) and presented the results in Table 2. To do the evaluation, we made use of the global test dataset to evaluate the global model performance following  (McMahan et al., 2017), and adopted the weighted average performance from client models to evaluate the plp following  (Fallah et al., 2020).
@@ 5.4 Overall Performance
MMiC and its two variants secure the top positions across all tasks and all missing modality scenarios. While MMiC (MR) excelled in classification tasks, being the only algorithm to achieve an F1 score exceeding 30 on CrisisMMD (mm@0.2, mc@0.5) at gbp. MMiC (GF) significantly outperformed all other algorithms in modality retrieval tasks across all three metrics. These results highlighted the superiority of MMiC in handling diverse missing modality scenarios.
@@
FedOpt demonstrated strong gbp on the CrisisMMD dataset (mm@0.2, mc@0.3) and (mm@0.2, mc@0.5). FedSoft’s performance is medium across experiments with multiple missing modalities settings and datasets both in global and personalized evaluations. As a CFL algorithm, PACFL do not exhibit outstanding performance in gbp. In contrast, it demonstrated excellent performance in personalized evaluation. CreamFL is primarily an algorithm designed for retrieval tasks. The overall performance of CreamFL is relatively mediocre except for Flickr30k (mm@0.2, mc@0.5), their method ranked second in this setting both in global and personalized evaluations. PmcmFL currently supports only classification tasks; however, its performance in classification tasks is somewhat limited.
(a) only PPS
(b) w/o PPS
(c) only BPI
(d) w/o BPI
(e) only MPO
(f) w/o MPO
Figure 2: Ablation study under CrisisMMD (mm@0.5 mc@0.5).
5.5 Ablation Study
@@
We conducted ablation experiments of MMiC on CrisisMMD. The results are shown in Figure 2. A further ablation study on Food101 and Flickr30k is provided in Appendix B. BPI (b), MPO (p) and PPS (s) represent Banzhaf Power Selection, Markowitz Portfolio Optimization, and Poverty Parameter Substitution, respectively. The figures show that MMiC consistently achieves higher accuracy in the final rounds across multiple ablation tests. Specifically, both ‘w/o PPS’ (Figure 2(b)) and ‘w/o BPI’ (Figure 2(d)) saw a decrease of more than 5% compared with MMiC. The final performance of ‘w/o MPO’ (Figure 2(f)) is close to MMiC, yet it remained approximately 2% lower than MMiC. In addition, removing any of the two contributions leads to varying degrees of performance decline, with the resulting performance consistently falling between that of MMiC and the configuration with a single module removed (Figure 2(a), 2(c), 2(e)). We conclude that removing any module leads to performance degradation. We further proved the effectiveness of BPI in Appendix C.
5.6 Experiments on other modalities
@@
We further experimented on two additional multimodal datasets with modalities other than image and text, aiming to show that MMiC is modality-agnostic. UCI-HAR (Reyes-Ortiz et al., 2013) and KU-HAR (Sikder & Nahid, 2021) both encompass smartphone sensor data utilized for the classification of diverse human activities. They include accelerometer and gyroscope readings gathered from a multitude of participants. We compared MMiC with FedOpt, which exhibited the good performance in Section 5.4 , for this analysis. The experimental setup involved 30 clients running 150 global rounds, with all other parameters remaining consistent with those outlined in Section 5.3. The findings of this experiment are presented in Table 3. MMiC demonstrated a pronounced superiority over FedOpt, with enhancements of 2% and 6% in the KU-HAR and UCI-HAR respectively.
Table 3: Performance comparison of MMiC and FedOpt on UCI-HAR and KU-HAR datasets.
@@
UCI-HAR 86.99 80.89
KU-HAR 75.83 73.24
6 Conclusions
Appendix A Learning Settings
In all of our experiments, we set a global learning rate (0.1, 0.01, 0.05) in the three datasets (CrisisMMD, Food101, Flickr30k) respectively to global optimizers ($FedAdagrad_{MPO}$, $FedAdagrad$). Our Parameter difference calculation process is derived from Equation 13 and we used the Adam optimizer with a learning rate of 0.005 for the clients. These parameters were obtained through our simple tuning in the early stage of our experiment. The batch size for the CrisisMMD dataset was set to 15, and 7 for the UPMC-Food101 dataset, 20 for the Flickr30k dataset. All the local epoch was set to 3. The loss function on the client was cross-entropy, except for MMiC (GMC-Fusion), where the loss consists of two parts: cross-entropy and contrastive loss between existing modalities (Poklukar et al., 2022).
Appendix B Ablation Study on Food101 and Flickr30k
We conducted an ablation study on the Food101 and Flickr30k datasets under the settings of mm@0.2 and mc@0.5. The modules are denoted as ‘b’, ‘p’, and ‘s’ (BPI, MPO and PPS respectively following Section 5.5). The full MMiC framework demonstrated robust performance, achieving an accuracy exceeding 90% on Food101 and a score of 11 on Flickr30k under the specified conditions. In contrast, the exclusion of all modules lead to a significant decline in accuracy, reducing it to 86% on Food101 and 4.3 on Flickr30k. The removal of a single module generally yields superior performance compared to the exclusion of two modules, as presented in Table 4.
Table 4: Ablation study on Food101 and Flickr30k under (mm@0.2, mc@0.5).
Configuration Food101 (Accuracy) Flickr30k (RSum)
w/o (b, p, s) 86.97 4.28
w/o s 88.54 6.99
w/o p 88.91 8.18
w/o b 87.26 8.10
only s 88.03 7.08
only p 87.49 4.84
only b 87.41 5.36
MMiC 90.12 11.02
@@ Appendix B Ablation Study on Food101 and Flickr30k
Figure 3: The number of additional training rounds needed for the Random Client Selection to achieve the same performance as the Banzhaf Client Selection (Flickr30k).
Appendix C Effectiveness of Banzhaf Client Selection
@@
The federated architecture had shown the feasibility of client selection, as demonstrated in (Nishio & Yonetani, 2019; Cho et al., 2020). The Banzhaf client selection strategy aimed to optimize global aggregation efficiency by minimizing communication rounds. Figure 3 illustrates the comparison of communication rounds and performance between the setting with Banzhaf client selection and random client selection across three levels of missing modality. We used Flickr30k dataset in this evaluation. It should be noted that randomly selected patterns required additional (12, 24, or 16) training rounds to surpass or approach the performance of the Banzhaf client selection.
Appendix D Time Complexity of MPO

### v3

@@ 4.4. Variants of MMiC
5. Experiments
5.1. Experiment Setup
Table 1. Global and personalized performance compared with different algorithms. Bold with * indicated the best performance, while bold black indicated excellent or stable top three performance. Since the PmcmFL algorithm is only for classification tasks, it can’t be compared in Flick30k. MR stands for Modality Reconstruction and GF is GMC-Fusion.
@@ 5.1. Experiment Setup
5.1.1. Datasets, Models and Metrics
Our evaluation of MMiC is based on two multimodal classification datasets and one multimodal retrieval dataset. Because each dataset has distinct characteristics, we followed (Feng et al., 2023; Jeni et al., 2013) in employing different base models, learning configurations and evaluation metrics for each dataset.
CrisisMMD (Ofli et al., 2020). It is a multi-modal dataset for classification consisting of 18,126 samples annotated tweets and images. We brought in a lightweight multi-modal model that used pre-trained embedding layers for input and concatenates shared representations (Gupta et al., 2024; Feng et al., 2023). Following (Feng et al., 2023), we employed the F1 (Goutte and Gaussier, 2005) score to evaluate the performance of CrisisMMD.
UPMC-Food101 (Gallo et al., 2020). UPMC-Food101 sourced from Food101, containing 101,000 image-text pairs across 101 food categories. We used CLIP with the official ViT version of the backbone pre-trained model (Radford et al., 2021). We froze the first 11 layers of attention while retaining the last layer of attention and the final output linear layer for parameter updates. We employed accuracy to evaluate its performance.
@@ 5.1.1. Datasets, Models and Metrics
Learning settings. In all of our experiments, we set a global learning rate (0.1, 0.01, 0.05) in the three datasets (CrisisMMD, Food101, Flickr30k) respectively to global optimizers (both $FedAdagrad$ and $FedAdagrad_{MPO}$). Our Parameter difference calculation process is derived from Equation 6 and we used the Adam optimizer with a learning rate of 0.005 for the clients. These parameters were obtained through our simple tuning in the early stage of our experiment. The batch size for the CrisisMMD dataset was set to 15, and 7 for the UPMC-Food101 dataset, 20 for the Flickr30k dataset. All the local epoch was set to 3. The loss function on the client was cross-entropy, except for MMiC (GMC-Fusion), where the loss consists of two parts: cross-entropy and contrastive loss between existing modalities (Poklukar et al., 2022).
5.1.2. Benchmark Algorithms
We compared MMiC with selected representative works and state-of-the-art works. FedAvg (McMahan et al., 2017) (first work for FL), FedOpt (Reddi et al., 2021b) (outstanding FL work), FedSoft (Ruan and Joe-Wong, 2022) (a soft SOTA clustered FL), PACFL (Vahidian et al., 2023) (an advanced CFL methodology), CreamFL (Yu et al., 2023) (MFL for addressing missing modality), and PmcmFL (Bao et al., 2024) (a SOTA work for addressing missing modality).
5.1.3. Cluster and Clients Settings
@@
In this experiment, we employed two distance metrics for clustering: Locality Sensitive Hashing (LSH) sketch (Yang et al., 2024; Liu et al., 2023) and SVD decomposition (Vahidian et al., 2023). All experiments used the LSH sketch for clustering unless otherwise noted. Each cluster had a test set visible to all clients within the cluster. Additionally, the global server provided accessibility to the test set for all clients. We simulated the Non-IID scenario by utilizing the Dirichlet distribution (Teh, 2010) with $\beta_{diri}=0.5$. For all the experiments, we set the number of clients to 50 so that ensured each client had at least 200 training samples. Our client selection strategy comprised an initial phase of 100 rounds with random selection, followed by a subsequent phase of 200 rounds utilizing Banzhaf client selection (300 global rounds).
5.1.4. Missing Modality Settings.
@@
Our setting with missing modalities was governed by two variables: clients with missing modality (mc) and missing modality rate (mm), denoting the proportion of all clients selected to have missing modalities and the proportion of missing samples on these clients, respectively. Our main experiment (Section 5.2) offered three levels of modality incompleteness settings: (mm@0.2, mc@0.3), (mm@0.2, mc@0.5) and (mm@0.5, mc@0.5). (mm@0.2, mc@0.3) means ratio of mm = 20% and ratio of mc = 30%. Each client has a fixed number of missing samples, determined by the mm rate. In each round, among the clients selected for aggregation, we use mc to decide whether the selected clients will activate missing data. Since the selected clients use a random training subset for training in each round, the maximum missing modality rate for a client in that round is $min(\frac{missing\ samples}{subset\ samples},1)$.
5.2. Overall Performance
@@
We evaluated MMiC based on global performance (gbp) and personalized performance (plp) and presented the results in Table 1. To do the evaluation, we made use of the global test dataset to evaluate the global model performance following  (McMahan et al., 2017), and adopted the weighted average performance from client models to evaluate the plp following  (Fallah et al., 2020). MMiC and its two variants secure the top positions across all tasks and all missing modality scenarios. While MMiC (MR) excelled in classification tasks, being the only algorithm to achieve an F1 score exceeding 30 on CrisisMMD (mm@0.2, mc@0.5) at gbp. MMiC (GF) significantly outperformed all other algorithms in modality retrieval tasks across all three metrics. These results highlighted the superiority of MMiC in handling diverse missing modality scenarios.
@@
FedOpt demonstrated strong gbp on the CrisisMMD dataset (mm@0.2, mc@0.3) and (mm@0.2, mc@0.5). FedSoft’s performance is medium across experiments with multiple missing modalities settings and datasets both in global and personalized evaluations. As a CFL algorithm, PACFL do not exhibit outstanding performance in gbp. In contrast, it demonstrated excellent performance in personalized evaluation. CreamFL is primarily an algorithm designed for retrieval tasks. The overall performance of CreamFL is relatively mediocre except for Flickr30k (mm@0.2, mc@0.5), their method ranked second in this setting both in global and personalized evaluations. PmcmFL currently supports only classification tasks; however, its performance in classification tasks is somewhat limited. Overall, MMiC consistently outperforms traditional methods for addressing missing modality challenges in FL.
5.3. Ablation Study
@@
We conducted an ablation study on the CrisisMMD, Food101, and Flickr30k datasets under the settings of mm@0.2 and mc@0.5 (Table 2). The ablation modules are denoted as ‘b’, ‘p’, and ‘s’ (BPI, MPO and PPS respectively). It shows that MMiC consistently achieves higher accuracy in the final rounds across multiple ablation tests. Specifically, ‘w/o b’ and ‘w/o s’ saw a decrease of more than 5% compared to MMiC in CrisisMMD while others still performed poorly. The final performance of ‘w/o p’ is close to MMiC, yet it remained approximately 2%, 0.2% and 3% lower than MMiC in three datasets respectively. In addition, removing any of the two contributions leads to varying degrees of performance decline, with the resulting performance consistently falling between that of MMiC and the configuration with a single module removed in Food101 and Flickr30k. We conclude that removing any module leads to performance degradation. We further proved the effectiveness of BPI in Section 5.5 and MPO in Section 5.6. We believe that BPI helps MMiC by reducing the weight of clients heavily affected by missing modalities, ensuring well-performing clients are prioritized. MPO further improves aggregation by minimizing the influence of high-risk clusters with missing modalities.
Table 2. Ablation study on CrisisMMD, Food101 and Flickr30k under (mm@0.2, mc@0.5).
Configuration CrisisMMD (Acc) Food101 (Acc) Flickr30k (RSum)
w/o (b, p, s) 57.22 86.97 4.28
w/o s 58.91 88.54 6.99
w/o p 61.07 88.91 8.18
w/o b 57.03 87.26 8.10
only s 60.36 88.03 7.08
only p 59.87 87.49 4.84
only b 61.24 87.41 5.36
MMiC 63.55 90.12 11.02
5.4. Experiments on other modalities
@@
In this section, we further broaden our experimental evaluation by incorporating more modalities, aiming to investigate the feasibility of our algorithm on non-dominant modalities, aiming to show that MMiC is modality-agnostic. UCI-HAR (Rasekh et al., 2014) and KU-HAR (Sikder and Nahid, 2021) both encompass smartphone sensor data utilized for the classification of diverse human activities. They include accelerometer and gyroscope readings gathered from a multitude of participants. We compared MMiC with FedOpt, which exhibited the good performance in Section 5.2 , for this analysis. The experimental setup involved 30 clients running 150 global rounds, with all other parameters remaining consistent with those outlined in Section 5.1.3. The findings of this experiment are presented in Table 3. MMiC demonstrated a pronounced superiority over FedOpt, with enhancements of 2% and 6% in the KU-HAR and UCI-HAR respectively.
Table 3. Performance comparison of MMiC and FedOpt on UCI-HAR and KU-HAR datasets.
@@
UCI-HAR (Acc) 86.99 75.83
KU-HAR (Acc) 80.89 73.24
@@
Figure 2. The number of additional training rounds needed for the Random Client Selection to achieve the same performance as the Banzhaf Client Selection (Flickr30k). The blue bar means random selection, and the yellow represents the BPI selection.
5.5. Effectiveness of Banzhaf Client Selection
@@
The federated architecture had shown the feasibility of client selection, as demonstrated in (Nishio and Yonetani, 2019; Cho et al., 2020). In this section, we compared training efficiency under missing modalities and show that BPI achieves higher efficiency than random selection. The Banzhaf client selection strategy aimed to optimize global aggregation efficiency by minimizing communication rounds. Figure 2 illustrates the comparison of communication rounds and performance between the setting with Banzhaf client selection and random client selection across three levels of missing modality. We used Flickr30k dataset in this evaluation. It should be noted that randomly selected patterns required additional (12, 24, or 16) training rounds to surpass or approach the performance of the Banzhaf client selection. To some extent, MMiC facilitates faster convergence of the global model under missing modality settings by prioritizing the selection of core clients.
5.6. MPO: A Potent Solution against Over-fitting
In this section, we show additional features of MPO against overfitting. We only implemented MMiC with MPO and deactivated PPS and BPI. Therefore, the comparison is between $FedAdagrad_{MPO}$ (MMiC) and $FedAdagrad$ (FedOpt) on dataset CrisisMMD (mm@0.2, mc@0.3) . Our purpose is to evaluate how the dynamic updating of parameters in $FedAdagrad_{MPO}$ contributes to enhanced stability. During the training, it had been converged after 400 rounds. We continued the training after 400 rounds while observing the personalized performance of $FedAdagrad$ and $FedAdagrad_{MPO}$, which showed the stability of our algorithm. The personalized evaluation is based on the average weighted performance of each client. It can be observed that the personalized performance of FedOpt begins to decline after round 413 and never recovers to its peak level thereafter. This indicated that some clients have already experienced overfitting, leading to a degradation in performance. In contrast, although the performance curve of MMiC experiences a short decline after round 543 as shown in Figure 3. $FedAdagrad_{MPO}$ forms risk by calculating the covariance of historical multi-round returns between pairs of clients (Equation 12). When partial clients overfit, the covariance increases, leading to a higher risk and an overall increase in RARC. When the majority of clients overfit, the total return decreases, yet the overall RARC still increases. An increase in RARC indicates that fewer parameters will be aggregated to the global model. The experimental results demonstrated that the stability of our algorithm is significantly enhanced by incorporating the MPO, particularly in mitigating performance degradation caused by overfitting during the training of clients.
Figure 3. Personalized performance of comparison after continuous training since the model convergence (400 rounds). The two dotted lines are the average scores. The upper part (blue) represents $FedAdagrad_{MPO}$ (MMiC) and the lower part (yellow) represents FedAdagradFedAdagrad (FedOpt).
Figure 4. The two plots on the left depict the SCV of LSH clustering under two different hyperparameter (number of clusters), while the two plots on the right show the SCV of SVD-based clustering with the threshold hyperparameter. The numbers within the images represent the indices of the cluster labels.
5.7. Global Performance Influenced by Clustered Federated Algorithms
We explored the impact of clustering quality on MMiC performance. Figure 4 plots Silhouette Coefficient Value (SCV) of two clustering algoritms with two different hyperparameters respectively. This experiment was conducted on dataset CrisisMMD (mm@0.2, mc@0.5) and we investigated how to change clustering hyperparameters to get a higher SCV. The SCV quantifies the ratio of the intra-cluster distance, defined as the distance from the sample point to the nearest cluster to which it doesn’t belong (Shahapure and Nicholas, 2020). A larger SCV signifies a greater separation between the cluster structure containing the sample point and its nearest neighbouring cluster structure, thereby indicating superior clustering quality under these conditions. For Locality Sensitive Hashing (LSH) clustering, their clustering algorithm is based on KMeans (Kanungo et al., 2002), which requires specifying the number of clusters. We adjusted the number of clusters from 1 to 20 to identify the two highest SCV. When the number of clusters is set to 3, the SCV attained its peak at 0.185. With 7 clusters and its second highest value of 0.125. For Singular Value Decomposition (SVD) clustering, which employs hierarchical clustering (Johnson, 1967), we varied the threshold from 1 to 20 to determine the optimal thresholds, identifying two values: 10 (SCV = 0.18) and 14 (SCV = 0.2). This preliminary selection of clustering parameters assisted us in identifying better and more suitable clustering algorithms and parameters for MMiC.
Figure 5 investigated the selection process of our clustering strategies. What we found are: (1) It’s shown that Poverty Parameter Substitution within clusters heavily relies on the quality of client clustering. For instance, in LSH clustering, a higher clustering quality (a higher SCV) often correlates with better global performance (LSH: 0.185 compared with LSH: 0.126). (2) By comparing different clustering methods, we can infer the contribution of each clustering algorithm to MMiC based on global performance. It is evident that LSH clustering outperformed SVD 0.2 in terms of global model performance when the SCV reached at 0.185. (3) We included manual clustering as a gold standard (clustering based on the overlap of client labels). Overall, both LSH (0.185) and SVD (0.2) clustering achieved global model performance comparable to that of manual clustering, making them reliable references for the clustering algorithms employed in this study. When compared to SVD, LSH demonstrated slightly superior performance.
Figure 5. We compared two clustering algorithms, LSH and SVD, by selecting the top-2 SCV hyperparameters for each and evaluating their global performance on MMiC. The blue curve is manual clustering.
5.8. Time Complexity of MPO
@@
6. Conclusions
7. GenAI Usage Disclosure
The authors acknowledge the use of an AI language model to assist in the linguistic refinement of selected sections of this paper, specifically the Introduction and Methodology. All the wordings, equations, and conclusions were revised again after being reviewed by the authors.